OpenStudy (anonymous):

Simple Calc 1 question, having a block for some reason. question and work in second post.

OpenStudy (anonymous):

\[f(x)=x^2(x-2)^4\] find the derivative use product rule \[2x(x-2)^4+x^2(4(x-2)^3(1))\] this is silly but this is the step i'm stuck on. I remember something about the x^2 but I cant remember why I take it out or what to do with it. any hints. please don't provide solutions. I just wanna get past my little bump so I can solve it

OpenStudy (shamil98):

Product and chain rule, i believe is required.

OpenStudy (anonymous):

shamils got it '-'. hes alreadt here

OpenStudy (shamil98):

Do you want to re-attempt it? Or should I give the step-by-step derivative?

OpenStudy (anonymous):

talking to me or lonnie?

OpenStudy (anonymous):

well I don't want the answer. but I can apply the chain rule to the first part. but what do I do with the x^2

OpenStudy (shamil98):

\[x^2 (x-2)^4\] \[2x(x-2)^4 + 4x^2(x-2)^3*1\] well actually you've done it correctly now that i've worked it out..

OpenStudy (anonymous):

when I apply the chain rule to the first part I get \[8(x-2)^3(1)+x^2(4(x-2)^3)\] something about the x^2 is making me wonder about something.. I cant remember.

OpenStudy (shamil98):

The product and chain rule combined is: \[\large f(x) * g(h(x))\] \[\large f'(x)*g(h(x)) + f(x) * g'(h(x))*h'(x)\] You did it correctly, nice job.

OpenStudy (shamil98):

Where did you get the 8 from?

OpenStudy (shamil98):

all you do to x^2(4(x-2)^3)) is multiply the 4 to x^2 4x^2(x-2)^3

OpenStudy (anonymous):

there has to be another step the solution shows to be \[2x(x-2)^3(3x-2)\] just ignore that 8. its not correct that's just me screwing around

OpenStudy (shamil98):

\[f(x) = x^2 \] \[g(h(x)) =(x-2)^4\] \[f'(x) = 2x\] \[g'(h(x))= 4(x-2)^3\] \[h(x) = 1\]

OpenStudy (shamil98):

Wolfram makes mistakes sometimes..

OpenStudy (shamil98):

\[2x(x-2)^4 + x^2*4(x-2)^3 * 1\] \[2x(x-2)^4 + 4x^2(x-2)^3\]

OpenStudy (shamil98):

h'(x) = 1 ** small typo there.

OpenStudy (anonymous):

it was wolfram and the book. is there a rule that requires to pull the x^2 out in front of the derivative or something? I see it in my book but I don't understand why its doing it if you factor out the 2x there you get \[(2x)((x-2)^4+2(x-2)^3)\]

OpenStudy (shamil98):

you made a typo i assume it would be \[(2x)((x-2)^4 + 2x(x-2)^3)\]

OpenStudy (shamil98):

And no there is no rule for that.

OpenStudy (anonymous):

ya my bad. now that makes it close to the book answer.

OpenStudy (shamil98):

I think the book is wrong, because using the product and chain rule combined, you should arrive at the answer... \[2x(x-2)^4 + 4x^2(x-2)^3\]

OpenStudy (shamil98):

No idea how it got this xD \[2x(x-2)^3(3x-2)\]

OpenStudy (anonymous):

Alright, I know the book could be wrong. I'm just a little obsessive about problems I can't solve. I'll leave it at that, thanks for the help. Yea I don't see where it came from either.

OpenStudy (shamil98):

Oh, never mind, the solution is the same thing!

OpenStudy (shamil98):

\[2x(x-2)^4 + 4x^2(x-2)^3 = 2x(x-2)^3(3x-2)\]

OpenStudy (shamil98):

I typed in the solution i got into wolfram, it's alternate form is the one from your book.

OpenStudy (shamil98):

Although I'm still skeptical as to how it is the same thing lol.

OpenStudy (anonymous):

is there a way to prove those are equal somehow lol

OpenStudy (shamil98):

We could solve it i guess.

OpenStudy (anonymous):

I guess I could use inductive reasoning and plug a 2 in there. that's a nono though

OpenStudy (shamil98):

\[2x((x-2)^4 +2x(x-2)^3) = 2x((x-2)^3(3x-2)\] dividing both sides by 2x \[(x-2)^4 + 2x(x-2)^3 = (x-2)^3(3x-2)\]

OpenStudy (shamil98):

From there i guess you would expand it and simplify.

OpenStudy (anonymous):

I plugged a 1 in for x and got 0=-2

OpenStudy (shamil98):

I would expand it but it would be a huge huge equation..

OpenStudy (shamil98):

\[(x-2)(x-2)(x-2)(x-2) + 2x(x-2)(x-2)(x-2)=(x-2)(x-2)(x-2)(3x-2)\] and then you would multiply i guess.

OpenStudy (anonymous):

same here lol. if I plugged in 1 and it didn't work it cant be true. if you plug 0 in only . its true. does that mean anything to you? couldn't we just start marking out x-2's on each side

OpenStudy (shamil98):

We can't we have +2x(x-2)^3

OpenStudy (shamil98):

we could have if it was multiplied by that ..

OpenStudy (anonymous):

\[2x(x-2)^4 + 4x^2(x-2)^3\] \[=(x-2)^3(2x(x-2)+4x^2)\]

OpenStudy (anonymous):

both terms have a common factor of \((x-2)^3\)

OpenStudy (shamil98):

Oh wait, yes we can, i'm not thinking straight..

OpenStudy (anonymous):

the second factor is \(4x^2-4x\) so maybe books answer is \[(x-2)^3(4x^2-4x)\] or even \[4(x-2)^3(x^2-x)\]

OpenStudy (shamil98):

\[(x-2)^3 ((x-2)+2x) = (x-2)^3(3x-2)\] \[(x-2) + 2x = 3x-2\] \[3x - 2 = 3x -2\] so yes it is correct!

OpenStudy (shamil98):

\[(x-2)^4 + 2x(x-2)^3 = (x-2)^3(3x-2)\] As sattelite said the left side has a gcf of (x-2)^3 which is why it is \[(x-2)^3 ((x-2)+2x)\]

OpenStudy (anonymous):

there must be many variations of this answer. Just need to have more faith in my answer. would you guys say that one of them is more 'correct' than the other?

OpenStudy (shamil98):

I think the derivative that I got is the best, because you would have to simplify it a lot more to get the other way.

OpenStudy (shamil98):

But, your book gave that simplified answer so next time i guess you are supposed to simplify it all the way

OpenStudy (anonymous):

possibly. thanks for the help. both of you. 3 different angles really helps

OpenStudy (shamil98):

Anytime!